If 6 people can cut 3820 square feet of shrubs in 1 hour, how many square feet can be cut by 8 people in 2 hours?
Answer by stanbon(75887)

If 6 people can cut 3820 square feet of shrubs in 1 hour, how many square feet can be cut by 8 people in 2 hours?
----
# of squae feet is directly related to # of people and # of hours.
----
Area = k*p*t
----
Solve for "k" using "6 people can cut 3820 square feet of shrubs in 1 hour".
----
3820 = k*6*1
k = 636 2/3
----
Equation:
A = (636 2/3)*p*t
----
how many square feet can be cut by 8 people in 2 hours?
A = (626.6667)*8*2
A = 10186 2/3 sq. ft.
=======================
Cheers,
Stan H.
===============